# A Cross-Domain Benchmark for Active Learning

### v1

@@ 1 Introduction
Deep neural networks (NN) have produced state-of-the-art results on many important supervised learning tasks. Since Deep NNs usually require large amounts of labeled training data, Active Learning (AL) deals with selecting the most informative samples out of a large pool of unlabeled data, so that only these samples need to be labeled. It has been shown that a small labeled set of this nature can be used to train well-performing models. In the last decade, many different algorithms for AL have been proposed and almost every method has reported lifts over all its predecessors.  [^1] Out of all considered algorithms for this paper, only BALD [8] did not claim a new state-of-the-art (SOTA) performance in their result section. However, real insights into the current state of AL are hard to draw from these works, due to the following reasons: 1. These works do not use a standardized evaluation setting with fixed datasets and baseline approaches. 2. Due to computational constraints, a lot of works perform only a small amount of experimental runs, hence it is questionable wether the superiority of a specific approach can be concluded from the conducted experiments. 3. The works are only evaluated in a specific domain, such as computer vision or language processing. However, AL is a general principle of supervised learning, and thus methods should be evaluated in multiple domains to assess their capabilities.
@@
Zhan et al. [29] sngl+batch 35 18 - - - 10-100
@@
Ours sngl+batch 9(14) 11 50
@@
To enable the computation of an oracle performance for a protocol with large amounts of restarts, we propose a greedy oracle algorithm which uses only a small amount of search steps to estimate the optimal solution. While being more time-efficient than established oracle functions, it possibly underestimates the real upper bound performance. However, as our experiments will show, it is still outperforming all current AL methods by at least 5% and thus is suitable as an upper bound.
@@
In summary, CDALBench is an experimental framework which includes an efficient oracle approximation, multiple application domains, enough repetitions to draw valid conclusions and two synthetic tasks to highlight shortcomings of AL methods. By being the first benchmark to providing these points in one code-base, we believe that CDALBench is a major step forward of assessing the overall state of AL research, independent of specific application domains. CDALBench is publicly available under https://anonymous.4open.science/r/A-Cross-Domain-Benchmark-for-Active-Learning-CCB1/README.md Our contributions include the following:
@@ 5 CDALBench: A Cross-Domain Active Learning Benchmark
Semi Splice Linear 100 o o o o
@@ 7.2 Results on Real-world Data
In Table 3 we provide the rank of each AL method per dataset. Please note, that we are averaging not only over runs, but also over query sizes per dataset, impacting AL methods that do not adapt well to a wide range of query sizes. For the results per query size, please refer to App. K. As mentioned in contribution 3, our results on real-world data show significant differences in the performance of methods between data domains: Not only do some methods overperform on some domains (like least confidence (LC) sampling on images), but the Top-3 of methods (except oracle) does not contain the same three methods for any two domains. Most interestingly, the image domain, which received most of the attention in benchmarking so far could even be considered an outlier, as this is the only domain where the Top-1 method changes. This highlights the dire need for diverse data domains in AL benchmarking. Results for the semi-supervised domain appear mostly in line with the other 3 domains, but a closer analysis of performances split into encoded images and encoded tabular reveals the need for further research. For details, please refer to App. L.
@@
Oracle 1.0 $\pm$ 0.01 1.0 $\pm$ 0.01 1.0 $\pm$ 0.0 1.0 $\pm$ 0.0 1.0 $\pm$ 0.0 1.0 $\pm$ 0.01 1.0 $\pm$ 0.0 1.0 2.0
Margin 6.6 $\pm$ 0.02 4.3 $\pm$ 0.01 2.1 $\pm$ 0.01 6.3 $\pm$ 0.01 4.4 $\pm$ 0.0 2.4 $\pm$ 0.01 3.7 $\pm$ 0.0 4.3 4.2
Badge 5.2 $\pm$ 0.01 6.3 $\pm$ 0.01 2.9 $\pm$ 0.01 5.2 $\pm$ 0.01 4.7 $\pm$ 0.0 3.3 $\pm$ 0.01 3.5 $\pm$ 0.0 4.5 5.4
LeastConf 9.2 $\pm$ 0.02 10.3 $\pm$ 0.02 8.1 $\pm$ 0.02 2.1 $\pm$ 0.01 4.0 $\pm$ 0.0 7.9 $\pm$ 0.02 3.0 $\pm$ 0.01 6.4 6.5
DSA 7.4 $\pm$ 0.02 7.3 $\pm$ 0.01 7.5 $\pm$ 0.01 5.4 $\pm$ 0.01 5.1 $\pm$ 0.0 6.0 $\pm$ 0.02 7.3 $\pm$ 0.01 6.6 6.7
BALD 4.0 $\pm$ 0.01 4.7 $\pm$ 0.01 5.4 $\pm$ 0.01 12.0 $\pm$ 0.01 7.6 $\pm$ 0.0 7.6 $\pm$ 0.02 5.0 $\pm$ 0.0 6.6 7.6
CoreGCN 6.9 $\pm$ 0.01 4.9 $\pm$ 0.01 10.4 $\pm$ 0.01 7.6 $\pm$ 0.01 6.5 $\pm$ 0.01 4.0 $\pm$ 0.01 6.8 $\pm$ 0.0 6.7 8.2
Entropy 6.6 $\pm$ 0.02 3.9 $\pm$ 0.01 7.6 $\pm$ 0.01 7.6 $\pm$ 0.01 4.9 $\pm$ 0.01 9.8 $\pm$ 0.02 9.6 $\pm$ 0.0 7.1 6.5
LSA 6.1 $\pm$ 0.01 6.8 $\pm$ 0.01 5.3 $\pm$ 0.01 7.7 $\pm$ 0.01 10.6 $\pm$ 0.01 7.5 $\pm$ 0.01 7.3 $\pm$ 0.01 7.3 7.5
Random 9.0 $\pm$ 0.01 9.3 $\pm$ 0.01 5.3 $\pm$ 0.01 8.4 $\pm$ 0.01 11.1 $\pm$ 0.0 7.9 $\pm$ 0.01 8.0 $\pm$ 0.0 8.4 6.9
Coreset 7.1 $\pm$ 0.01 9.0 $\pm$ 0.01 10.5 $\pm$ 0.01 6.8 $\pm$ 0.01 7.1 $\pm$ 0.0 8.5 $\pm$ 0.02 10.8 $\pm$ 0.01 8.5 7.2
TypiClust 8.8 $\pm$ 0.01 10.2 $\pm$ 0.01 12.0 $\pm$ 0.02 7.9 $\pm$ 0.01 11.0 $\pm$ 0.01 12.0 $\pm$ 0.02 12.0 $\pm$ 0.01 10.5 9.2
@@ 8 Honeypot and Diverging Sine
9 Conclusion
@@ Appendix I Oracle Curve Forecasting
Unfortunately, the iterative nature of our oracle means that the computational effort scales in the budget $B$. For datasets with large budgets, like Cifar10 and FashionMnist (both 10K), we were unable to compute the oracle set for the entire 10K iterations. We compromised by (i) picking the two points with highest test accuracy, instead of only one and (ii) only computed until iteration 2000. The rest of the curve was forecast using a simple 2-step algorithm, based on linear regression:
@@
Uncertainty Sampling tries to find the sample that the classifier is most uncertain about by computing heuristics of the class probabilities. For our benchmark, we use entropy and margin (a.k.a. best-vs-second-best) sampling. BALD [14] applies the query-by-committee strategy of model ensembles to a single model by interpreting the classifier’s parameters as distributions and then sample multiple outputs from them via Monte-Carlo dropout. BADGE [1] uses gradient embeddings of unlabeled points to select samples where the classifier is expected to change a lot. The higher the magnitude of the gradient the higher the expected improvement in model performance. BADGE employs a variant to the KMeans++ initialization technique to select batches of points. Even though [1] provided pseudocode for this procedure that selects the first point at random, all found implementations of BADGE select the first points instead by maximum gradient magnitude. Coreset [25] employs K-Means clustering trying to cover the whole data distribution. Selects the unlabeled sample that is the furthest away from all cluster centers. Clustering is done in a semantically meaningful space by encoding the data with the current classifier $\hat{y}$. In this work, we use the greedy variant of Coreset. TypiClust [9] relies on clustering similar to Coreset, but proposes a new measure called “Typicality” to select unlabeled samples. It selects points that are in the densest regions of clusters that do not contain labeled samples yet. Clustering is done in a semantically meaningful space by encoding the data with the current classifier $\hat{y}$. It has to be pointed out that TypiClust was designed for low-budget scenarios, but we think it is still worthwhile to test and compare this method with higher budgets. Core-GCN [3] trains a Graph-Convolutional-Network (GCN) on embeddings of the unlabeled pool, obtained from the classifier (Similar to Coreset and TypiClust). This GCN model propagates uncertainty information through the graph and therefore enhances the nodes uncertainty quantification. Lastly, the node that displays the highest amount of uncertainty is selected for labeling. DSA/LSA [13] use the metric of test adequacy to construct a set of points that is diverse, ranging from points that are close to points in $\mathcal{L}$ and points that are significantly different. DSA and LSA measure the diversity of points by distance in embedding space or likelihood estimation under the given classifier respectively. Excluded Methods Learning Loss for AL [28] Introduces an updated training of the classification model with an auxiliary loss and therefore cannot be compared fairly against classification models without this boosted training regime. Reinforcement Learning Methods We postpone the study of learned AL methods to future versions of this benchmark, as reinforcement learning is infamous for being extremely time consuming and itself hard to reproduce .
@@
LeastConf 2 0.779 $\pm$ 0.019 0.68 $\pm$ 0.032 0.809 $\pm$ 0.01 0.629 $\pm$ 0.05 0.846 $\pm$ 0.009 0.421 $\pm$ 0.039 0.668 $\pm$ 0.014 0.685 $\pm$ 0.009 0.843 $\pm$ 0.013 0.94 $\pm$ 0.016 0.692 $\pm$ 0.094
DSA 3 0.766 $\pm$ 0.021 0.691 $\pm$ 0.022 0.803 $\pm$ 0.01 0.646 $\pm$ 0.032 0.829 $\pm$ 0.01 0.431 $\pm$ 0.05 0.663 $\pm$ 0.014 0.679 $\pm$ 0.01 0.844 $\pm$ 0.017 0.941 $\pm$ 0.014 0.731 $\pm$ 0.032
BALD 1 0.78 $\pm$ 0.014 0.649 $\pm$ 0.04 0.784 $\pm$ 0.01 0.632 $\pm$ 0.042 0.819 $\pm$ 0.01 0.242 $\pm$ 0.046 0.666 $\pm$ 0.014 0.644 $\pm$ 0.018 0.815 $\pm$ 0.024 0.928 $\pm$ 0.014 0.698 $\pm$ 0.043
CoreGCN 2 0.765 $\pm$ 0.021 0.686 $\pm$ 0.023 0.804 $\pm$ 0.012 0.646 $\pm$ 0.03 0.753 $\pm$ 0.016 0.39 $\pm$ 0.044 0.623 $\pm$ 0.018 0.647 $\pm$ 0.012 0.85 $\pm$ 0.01 0.938 $\pm$ 0.014 0.731 $\pm$ 0.028
Entropy 3 0.768 $\pm$ 0.022 0.678 $\pm$ 0.035 0.812 $\pm$ 0.013 0.635 $\pm$ 0.045 0.83 $\pm$ 0.011 0.399 $\pm$ 0.035 0.663 $\pm$ 0.013 0.681 $\pm$ 0.011 0.815 $\pm$ 0.021 0.942 $\pm$ 0.017 0.696 $\pm$ 0.083
@@
Margin 1 0.747 $\pm$ 0.023 0.751 $\pm$ 0.019 0.828 $\pm$ 0.009 0.363 $\pm$ 0.031 0.64 $\pm$ 0.013 0.653 $\pm$ 0.01 0.774 $\pm$ 0.029 0.46 $\pm$ 0.006
Badge 3 0.758 $\pm$ 0.017 0.754 $\pm$ 0.018 0.831 $\pm$ 0.008 0.376 $\pm$ 0.028 0.632 $\pm$ 0.013 0.649 $\pm$ 0.011 0.781 $\pm$ 0.026 0.462 $\pm$ 0.007
LeastConf 2 0.731 $\pm$ 0.025 0.688 $\pm$ 0.041 0.761 $\pm$ 0.037 0.291 $\pm$ 0.03 0.644 $\pm$ 0.013 0.65 $\pm$ 0.011 0.73 $\pm$ 0.049 0.462 $\pm$ 0.009
@@
Oracle 0 0.803 $\pm$ 0.012 0.825 $\pm$ 0.009 0.866 $\pm$ 0.004 0.436 $\pm$ 0.057 0.749 $\pm$ 0.009 0.755 $\pm$ 0.005 0.49 $\pm$ 0.003
Margin 2 0.733 $\pm$ 0.024 0.711 $\pm$ 0.027 0.799 $\pm$ 0.013 0.473 $\pm$ 0.026 0.629 $\pm$ 0.012 0.628 $\pm$ 0.009 0.455 $\pm$ 0.006
Badge 2 0.743 $\pm$ 0.014 0.714 $\pm$ 0.032 0.804 $\pm$ 0.013 0.472 $\pm$ 0.029 0.623 $\pm$ 0.01 0.621 $\pm$ 0.01 0.456 $\pm$ 0.006
LeastConf 0 0.715 $\pm$ 0.033 0.639 $\pm$ 0.05 0.708 $\pm$ 0.034 0.23 $\pm$ 0.034 0.631 $\pm$ 0.013 0.62 $\pm$ 0.012 0.457 $\pm$ 0.008
@@
Cifar10 FashionMnist
Oracle 0.689 $\pm$ 0.001 0.905 $\pm$ 0.001
Margin 0.556 $\pm$ 0.008 0.882 $\pm$ 0.004
Badge 0.56 $\pm$ 0.008 0.883 $\pm$ 0.005
LeastConf 0.591 $\pm$ 0.01 0.884 $\pm$ 0.005
DSA 0.56 $\pm$ 0.009 0.882 $\pm$ 0.004
BALD 0.478 $\pm$ 0.014 0.878 $\pm$ 0.003
CoreGCN 0.553 $\pm$ 0.01 0.88 $\pm$ 0.007
Entropy 0.553 $\pm$ 0.009 0.882 $\pm$ 0.006
LSA 0.558 $\pm$ 0.01 0.866 $\pm$ 0.005
Random 0.557 $\pm$ 0.01 0.863 $\pm$ 0.005
Coreset 0.553 $\pm$ 0.007 0.878 $\pm$ 0.006
TypiClust 0.557 $\pm$ 0.009 0.864 $\pm$ 0.004
@@
BADGE 100 Based on [1, 16]
BALD 100 Dropout Trials: 5 Based on [4]
Coreset 8000 Own
TypiClust 10000 Min Cluster Size: 5 Max # Clusters: 500 Based on [9]
Margin 8000 Own
Entropy 8000 Own

### v2

@@ 1 Introduction
Deep neural networks (NN) have produced state-of-the-art results on many important supervised learning tasks. Since Deep NNs usually require large amounts of labeled training data, Active Learning (AL) deals with selecting the most informative samples out of a large pool of unlabeled data, so that only these samples need to be labeled. It has been shown that a small labeled set of this nature can be used to train well-performing models. In the last decade, many different algorithms for AL have been proposed and almost every method has reported lifts over every single one of its predecessors.  [^1] Out of all considered algorithms for this paper, only BALD [9] did not claim a new state-of-the-art (SOTA) performance in their result section. However, real insights into the current state of AL are hard to draw from these works, due to the following reasons: 1. These works do not use a standardized evaluation setting with fixed datasets and baseline approaches. 2. Due to computational constraints, a lot of works perform only a small amount of experimental runs, hence it is questionable wether the superiority of a specific approach can be concluded from the conducted experiments. 3. The works are only evaluated in a specific domain, such as computer vision or language processing. However, AL is a general principle of supervised learning, and thus methods should be evaluated in multiple domains to assess their capabilities.
@@
Zhan et al. [30] single+batch 35 18 - - - 10-100
@@
Ours single+batch 9(14) 11 50
@@
To enable the computation of an oracle performance for a protocol with large amounts of restarts, we propose a greedy oracle algorithm which uses only a small amount of search steps to estimate the optimal solution. Our oracle relies on directly testing a small sample of points in every iteration whether they induce an improvement in test accuracy and selects the optimal point from that small sample. While being more time-efficient than established oracle functions, it possibly underestimates the real upper bound performance. However, as our experiments will show, it is still outperforming all current AL methods by at least 5% and thus is suitable as an upper bound.
@@
In summary, CDALBench is an experimental framework which includes an efficient oracle approximation, multiple application domains, enough repetitions to draw valid conclusions and two synthetic tasks to highlight shortcomings of AL methods. By being the first benchmark to providing these points in one code-base, we believe that CDALBench is a major step forward of assessing the overall state of AL research, independent of specific application domains. CDALBench is publicly available under https://github.com/wernerth94/A-Cross-Domain-Benchmark-for-Active-Learning/. Our contributions include the following:
@@ 5 CDALBench: A Cross-Domain Active Learning Benchmark
Semi Splice Linear 100 o o o
@@ 7.2 Results on Real-world Data
In Table 3 we provide the rank of each AL method per dataset. Please note, that we are averaging not only over runs, but also over query sizes per dataset, impacting AL methods that do not adapt well to a wide range of query sizes. For the results per query size, please refer to App. K. As mentioned in contribution 3, our results on real-world data show significant differences in the performance of methods between data domains: Not only do some methods overperform on some domains (like least confidence (LC) sampling on images), but the Top-3 of methods (except oracle) does not contain the same three methods for any two domains. Most interestingly, the image domain, which received most of the attention in benchmarking so far could even be considered an outlier, as this is the only domain where the Top-1 method changes. This highlights the dire need for diverse data domains in AL benchmarking. Results for the semi-supervised domain appear mostly in line with the other 3 domains, but a closer analysis of performances split into encoded images and encoded tabular reveals the need for further research. For details, please refer to App. L. Finally, we would like to emphasize that the total average rank of our top 3 algorithms (column ”Normal” in Tab. 3) are 4.8, 4.9 and 5.1. No single algorithm was able to perform well in every domain, either being outperformed by a specialist algorithm in each domain, or experiencing a severe drop in performance in a poorly matched domain.
@@
Oracle 1.0$\pm$0.01 1.0$\pm$0.01 1.0$\pm$0.0 1.0$\pm$0.0 1.0$\pm$0.0 1.0$\pm$0.01 1.0$\pm$0.0 1.0 2.1
Margin 6.8$\pm$0.02 4.5$\pm$0.01 2.7$\pm$0.01 7.2$\pm$0.01 4.9$\pm$0.0 3.1$\pm$0.01 4.5$\pm$0.0 4.8 4.7
Galaxy 9.5$\pm$0.02 9.4$\pm$0.02 2.4$\pm$0.01 2.7$\pm$0.01 5.6$\pm$0.01 2.6$\pm$0.01 2.0$\pm$0.0 4.9 5.7
Badge 6.2$\pm$0.01 6.5$\pm$0.01 3.8$\pm$0.01 6.2$\pm$0.01 5.1$\pm$0.0 4.2$\pm$0.01 3.6$\pm$0.0 5.1 6.0
LeastConfident 9.6$\pm$0.02 11.1$\pm$0.02 9.1$\pm$0.02 2.6$\pm$0.01 4.4$\pm$0.0 8.9$\pm$0.02 5.2$\pm$0.01 7.3 7.1
DSA 7.8$\pm$0.02 7.7$\pm$0.01 8.5$\pm$0.01 6.4$\pm$0.01 5.6$\pm$0.0 7.0$\pm$0.02 8.1$\pm$0.01 7.3 7.3
CoreGCN 7.2$\pm$0.01 5.2$\pm$0.01 11.4$\pm$0.01 8.6$\pm$0.01 7.1$\pm$0.01 5.0$\pm$0.01 7.4$\pm$0.01 7.4 8.9
BALD 4.1$\pm$0.01 4.8$\pm$0.01 6.4$\pm$0.01 13.0$\pm$0.01 8.4$\pm$0.0 8.6$\pm$0.02 7.7$\pm$0.0 7.6 8.3
Entropy 6.8$\pm$0.02 4.0$\pm$0.01 8.6$\pm$0.01 8.6$\pm$0.01 5.4$\pm$0.01 10.8$\pm$0.02 11.1$\pm$0.01 7.9 7.3
LSA 6.2$\pm$0.01 7.2$\pm$0.01 6.3$\pm$0.01 8.6$\pm$0.01 11.6$\pm$0.01 8.5$\pm$0.01 7.7$\pm$0.0 8.0 8.1
Random 9.4$\pm$0.01 9.7$\pm$0.01 6.3$\pm$0.01 9.4$\pm$0.01 12.1$\pm$0.0 8.9$\pm$0.01 8.3$\pm$0.0 9.2 7.6
Coreset 7.3$\pm$0.01 9.5$\pm$0.01 11.5$\pm$0.01 7.7$\pm$0.01 7.8$\pm$0.0 9.5$\pm$0.02 11.5$\pm$0.01 9.3 7.9
TypiClust 9.1$\pm$0.01 10.6$\pm$0.01 13.0$\pm$0.02 8.9$\pm$0.01 12.0$\pm$0.01 13.0$\pm$0.02 13.0$\pm$0.01 11.4 9.9
@@ 8 Honeypot and Diverging Sine
9 Comparison to other Benchmarks
Comparing our results to the findings of other works based in accuracy scores would be meaningless, as every work employs different models, hyperparameters and training loops. We instead opt to compare only the ranking of algorithms to the literature. Our results generally are reflected in domain-specific benchmarks - [32] also find least confidence sampling and BADGE to be the best algorithms for images (they don’t test Galaxy), [25] also find BADGE to be the best algorithm for text (they don’t test margin sampling or Galaxy) and [2] also find margin sampling to be the best algorithm for tabular data. In this work, we provide the first AL benchmark that spans all major data domains and is easily reproducible while matching and extending previous published results from single-domain benchmarks.
10 Using this Benchmark
@@ Appendix I Oracle Curve Forecasting
Unfortunately, the iterative nature of our oracle means that the computational effort scales in the budget $O(\tau B)$. For datasets with large budgets, like Cifar10 and FashionMnist (both 10K), we were unable to compute the oracle set for the entire 10K iterations. We compromised by (i) picking the two points with highest test accuracy, instead of only one and (ii) only computed until iteration 2000. The rest of the curve was forecast using a simple 2-step algorithm, based on linear regression:
@@
Uncertainty Sampling tries to find the sample that the classifier is most uncertain about by computing heuristics of the class probabilities. For our benchmark, we use entropy and margin (a.k.a. best-vs-second-best) sampling. BALD [15] applies the query-by-committee strategy of model ensembles to a single model by interpreting the classifier’s parameters as distributions and then sample multiple outputs from them via Monte-Carlo dropout. BADGE [1] uses gradient embeddings of unlabeled points to select samples where the classifier is expected to change a lot. The higher the magnitude of the gradient the higher the expected improvement in model performance. BADGE employs a variant to the KMeans++ initialization technique to select batches of points. Even though [1] provided pseudocode for this procedure that selects the first point at random, all found implementations of BADGE select the first points instead by maximum gradient magnitude. Coreset [26] employs K-Means clustering trying to cover the whole data distribution. Selects the unlabeled sample that is the furthest away from all cluster centers. Clustering is done in a semantically meaningful space by encoding the data with the current classifier $\hat{y}$. In this work, we use the greedy variant of Coreset. TypiClust [10] relies on clustering similar to Coreset, but proposes a new measure called “Typicality” to select unlabeled samples. It selects points that are in the densest regions of clusters that do not contain labeled samples yet. Clustering is done in a semantically meaningful space by encoding the data with the current classifier $\hat{y}$. It has to be pointed out that TypiClust was designed for low-budget scenarios, but we think it is still worthwhile to test and compare this method with higher budgets. Core-GCN [4] trains a Graph-Convolutional-Network (GCN) on embeddings of the unlabeled pool, obtained from the classifier (Similar to Coreset and TypiClust). This GCN model propagates uncertainty information through the graph and therefore enhances the nodes uncertainty quantification. Lastly, the node that displays the highest amount of uncertainty is selected for labeling. DSA/LSA [14] use the metric of test adequacy to construct a set of points that is diverse, ranging from points that are close to points in $\mathcal{L}$ and points that are significantly different. DSA and LSA measure the diversity of points by distance in embedding space or likelihood estimation under the given classifier respectively. GALAXY [33] construct a graph, where every point in $\mathcal{U}$ is a node and the edges are constructed from the current classifiers output. The algorithm then queries points from $\mathcal{U}$ based on edges in the graph, that are connected to two nodes with different predicted labels. Excluded Methods Learning Loss for AL [29] Introduces an updated training of the classification model with an auxiliary loss and therefore cannot be compared fairly against classification models without this boosted training regime. Reinforcement Learning Methods We postpone the study of learned AL methods to future versions of this benchmark, as reinforcement learning is infamous for being extremely time consuming and itself hard to reproduce .
@@
LeastConfident 3 0.779$\pm$0.019 0.68$\pm$0.032 0.809$\pm$0.01 0.629$\pm$0.05 0.846$\pm$0.009 0.421$\pm$0.039 0.668$\pm$0.014 0.685$\pm$0.009 0.843$\pm$0.013 0.94$\pm$0.016 0.692$\pm$0.094
DSA 4 0.766$\pm$0.021 0.691$\pm$0.022 0.803$\pm$0.01 0.646$\pm$0.032 0.829$\pm$0.01 0.431$\pm$0.05 0.663$\pm$0.014 0.679$\pm$0.01 0.844$\pm$0.017 0.941$\pm$0.014 0.731$\pm$0.032
BALD 0 0.78$\pm$0.014 0.649$\pm$0.04 0.784$\pm$0.01 0.632$\pm$0.042 0.819$\pm$0.01 0.242$\pm$0.046 0.666$\pm$0.014 0.644$\pm$0.018 0.815$\pm$0.024 0.928$\pm$0.014 0.698$\pm$0.043
CoreGCN 4 0.765$\pm$0.021 0.686$\pm$0.023 0.804$\pm$0.012 0.646$\pm$0.03 0.753$\pm$0.016 0.39$\pm$0.044 0.623$\pm$0.018 0.647$\pm$0.012 0.85$\pm$0.01 0.938$\pm$0.014 0.731$\pm$0.028
Entropy 2 0.768$\pm$0.022 0.678$\pm$0.035 0.812$\pm$0.013 0.635$\pm$0.045 0.83$\pm$0.011 0.399$\pm$0.035 0.663$\pm$0.013 0.681$\pm$0.011 0.815$\pm$0.021 0.942$\pm$0.017 0.696$\pm$0.083
@@
Margin 0 0.747$\pm$0.023 0.751$\pm$0.019 0.828$\pm$0.009 0.363$\pm$0.031 0.64$\pm$0.013 0.653$\pm$0.01 0.774$\pm$0.029 0.46$\pm$0.006
Galaxy 4 0.733$\pm$0.027 0.702$\pm$0.037 0.828$\pm$0.013 0.363$\pm$0.029 0.647$\pm$0.009 0.654$\pm$0.008 0.781$\pm$0.029 0.482$\pm$0.006
Badge 3 0.748$\pm$0.017 0.754$\pm$0.018 0.831$\pm$0.008 0.376$\pm$0.028 0.632$\pm$0.013 0.649$\pm$0.011 0.781$\pm$0.026 0.462$\pm$0.007
LeastConfident 0 0.731$\pm$0.025 0.688$\pm$0.041 0.761$\pm$0.037 0.291$\pm$0.03 0.644$\pm$0.013 0.65$\pm$0.011 0.73$\pm$0.049 0.462$\pm$0.009
@@
Oracle 0.803$\pm$0.012 0.825$\pm$0.009 0.866$\pm$0.004 0.436$\pm$0.057 0.749$\pm$0.009 0.755$\pm$0.005 0.49$\pm$0.003
Margin 1 0.733$\pm$0.024 0.711$\pm$0.027 0.799$\pm$0.013 0.473$\pm$0.026 0.629$\pm$0.012 0.628$\pm$0.009 0.455$\pm$0.006
Galaxy 4 0.715$\pm$0.032 0.658$\pm$0.047 0.804$\pm$0.015 0.307$\pm$0.04 0.638$\pm$0.01 0.632$\pm$0.009 0.478$\pm$0.007
Badge 1 0.743$\pm$0.014 0.714$\pm$0.032 0.804$\pm$0.013 0.472$\pm$0.029 0.623$\pm$0.01 0.621$\pm$0.01 0.456$\pm$0.006
LeastConfident 0 0.715$\pm$0.033 0.639$\pm$0.05 0.708$\pm$0.034 0.23$\pm$0.034 0.631$\pm$0.013 0.62$\pm$0.012 0.457$\pm$0.008
@@
Cifar10 FashionMnist News
Oracle 0.689$\pm$0.001 0.905$\pm$0.001 0.49$\pm$0.003
Margin 0.556$\pm$0.008 0.882$\pm$0.004 0.441$\pm$0.011
Galaxy 0.591$\pm$0.011 0.881$\pm$0.008 0.447$\pm$0.009
Badge 0.56$\pm$0.008 0.883$\pm$0.005 0.451$\pm$0.008
LeastConfident 0.591$\pm$0.01 0.884$\pm$0.005 0.425$\pm$0.013
DSA 0.56$\pm$0.009 0.882$\pm$0.004 0.432$\pm$0.014
BALD 0.478$\pm$0.014 0.878$\pm$0.003 0.421$\pm$0.012
CoreGCN 0.553$\pm$0.01 0.88$\pm$0.007 0.437$\pm$0.012
Entropy 0.553$\pm$0.009 0.882$\pm$0.006 0.419$\pm$0.015
LSA 0.558$\pm$0.01 0.866$\pm$0.005 0.437$\pm$0.009
Random 0.557$\pm$0.01 0.863$\pm$0.005 0.437$\pm$0.008
Coreset 0.553$\pm$0.007 0.878$\pm$0.006 0.439$\pm$0.009
TypiClust 0.557$\pm$0.009 0.864$\pm$0.004 0.43$\pm$0.01
@@
BADGE 5000 Based on [1, 17]
BALD 5000 Dropout Trials: 5 Based on [5]
Coreset 5000 Own
TypiClust 5000 Min Cluster Size: 5 Max # Clusters: 500 Based on [10]
Margin 5000 Own
Entropy 5000 Own